OpenStudy (anonymous):

Car A leaves Sydney for Brisbane at 8:00 am and traveled at an average speed of 60km/h. At 9:00am Car B also leaves Sydney for Brisbane. If Car B at an average speed of 80km/h, at what time will it overtake Car A?

OpenStudy (lgbasallote):

you do know you couldve just edited the question a while ago there and then right....anyway....

OpenStudy (anonymous):

No I didn't know that

OpenStudy (lgbasallote):

it's the link beside Report Abuse i thin..

OpenStudy (lgbasallote):

anyway... at 9:00 Car A = 60 km Car B = 0km 10:00 Car A = 120 km Car B = 80 km 11: 00 Car A = 180 Car B = 160 (getting near) 12:00 Car A = 240 Car B = 240 :O lookie! we have an overtaker yay!!! so...12:00 is how many hours from 9:00?

OpenStudy (lgbasallote):

uhhh @Genuine might wanna give some feedback there /:)

OpenStudy (anonymous):

@lgbasallote is right

OpenStudy (anonymous):

@lgbasallote Oh thanks! (:

OpenStudy (anonymous):

But is there like a way to solve this question with an equation?

OpenStudy (lgbasallote):

oh yeah...im thinking it out wait...

OpenStudy (lgbasallote):

i think i got it... let t = Car B's time thing (t+1) = Car A's time (because it starts 1 hr before Car B) d_A = d_B <---overtake 60(t+1) = 80(t) 60t + 60 = 80t 60 = 80t - 60t 60 = 20t 3 = t seems about right :DDD

OpenStudy (anonymous):

Why do people think I am you?

OpenStudy (anonymous):

@lgbasallote I never thought of that method (: But the example in my book says something about (t + 3/2) or something, would you know what it means?

OpenStudy (lgbasallote):

hmmm is it regarding this same problem? will you post the whole solution? maybe i can figure it out then

OpenStudy (anonymous):

No it's not the exact same question but it's very similar: Car A left Sydney for Melbourne at 6am and traveled at an average speed of 80km/h. At 7:30 am car B left Sydney for Melbourne. If car B travels at an average speed of 100km.h, at what time will it catch car A?

OpenStudy (lgbasallote):

ahhh that's where 3/2 came from...7:30 is 1 1/2 hrs from 6:00 aka 3/2 hrs

OpenStudy (anonymous):

And the solution for it is: Car B will catch A when both have traveled the same distance. Now D=ST and let time be T be the time that car B catches up with car A. For car A it travels for (T+3/2)hours and for a distance of 80(T+3/2) For car B it travels for T hours and for a distance of 100T. Thus the distance traveled by both cars is the same...

OpenStudy (lgbasallote):

....i already said where 3/2 came from @Genuine :P

OpenStudy (anonymous):

Oh yeah, thank you again for your help (:

OpenStudy (lgbasallote):

<tips hat>